Methyl benzoate, which has a molar mass of 136.15 g/mol, can be nitrated to methyl nitrobenzoate, which has a molar mass of 181.13 g/mol, by the addition of nitric acid.

What is the theoretical yield (in g) of the mononitrated product, methyl nitrobenzoate, from 0.560 g of methyl benzoate?
Answer #1

No. of moles of methyl benzoate = (0.560 g) / (136.15 g/mol) = 0.0041 mol

No. of moles of mononitrated product formed = (1 / 1) x (0.0041 mol) = 0.0041 mol

Theoretical yield of product produced = (0.0041 mol) x (181.13 g/mol) = 0.745 grams